OpenStudy (anonymous):

How do you factor -x^3+6x^2-11x+6 ??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is the problem that u r facing elaborate it

OpenStudy (anonymous):

i just need to solve it for 0.

OpenStudy (anonymous):

means that u want the RHS TO BE 0 or x be 0

OpenStudy (anonymous):

-x^3+6x^2-11x+6=0 and then solve for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

IV JUST WAIT FOR 2minutes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Putting x=1 we will get:

OpenStudy (anonymous):

-1 + 6 -11 + 6 = 0

OpenStudy (anonymous):

-12 + 12 = 0

OpenStudy (anonymous):

so it shows that x-1 is a factor of THE EQUATION

OpenStudy (anonymous):

I know the answers are 1 2 and 3. But how do factor this out is my question.

OpenStudy (anonymous):

(-x^3 + 6x^2 -11x + 6) * -1 = 0*-1

OpenStudy (anonymous):

ok now

OpenStudy (anonymous):

x^3 - 6x^2 + 11x - 6 = 0

OpenStudy (anonymous):

x^3 = x^2(x-1) +x^2

OpenStudy (anonymous):

-6x^2 = -6x(x-1) - 6x

OpenStudy (anonymous):

11x = 11 ( x-1) + 11

OpenStudy (anonymous):

putting these values in this equation x^2(x-1) - 5x(x-1) + 6(x-1) = 0

OpenStudy (anonymous):

(x^2-5x+)(x-1) = 0

OpenStudy (anonymous):

(x(x-2) - 3(x-2) ) (x-1) = 0

OpenStudy (anonymous):

x-3, x-2, x-1 = 0

OpenStudy (anonymous):

x-3 = 0 x=3 x-2 = 0 x=2 x-1 = 0 x=1

OpenStudy (anonymous):

SO BYE AND REPLY ME FAST did u got it

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

BYE